# Transitions to Intermittent Chaos in Quorum Sensing Dynamics

### v2

@@ 2 Nonlinear time delayed Quorum Sensing system
Figure 2: Purely imaginary eigenvalues for $\tau_{1},\tau_{2}>0$ for low and high frequency intervals. (a) Conditions in (11) hold for $\omega>0$ within the two distinct frequency intervals given by $\Omega_{1}=[0,0.0466554]$ and $\Omega_{2}=[0.136335,0.187117]$. (b)-(c) Sets of positive delay values for which the pseudo-polynomial (7) admits at least one zero of the form $\lambda=i\omega$ with $\omega\in\Omega_{1}=[0,0.0466554]$ (panel (b)) and $\omega\in\Omega_{2}=[0.136335,0.187117]$ (panel (c)). The asterisks mark points in (b) and (c) correspond to the specific case of $\tau_{1}=52$. Delay values satisfying the transversality parameter (13) to be positive are in red, otherwise correspond to blue solid curves. Other parameter values as in Table 1.
@@
Second, we consider the scenario where a pair of complex conjugate roots of $q(\lambda,\boldsymbol{\tau},\mathbf{p})$ crosses the imaginary axis into the right-hand side of the complex plane. In this case, the delay vector $\boldsymbol{\tau}=(\tau_{1},\tau_{2})$ shifts parallel to the vector $\mathbf{w}_{r}=(\partial\tau_{1}/\partial\sigma,\partial\tau_{2}/\partial\sigma)$. Consequently, if $\boldsymbol{\tau}\in\mathbb{R}_{+}^{2}$ lies to the left of a positively oriented trajectory $\mathcal{T}^{k}$, two additional roots of $q(\lambda,\boldsymbol{\tau},\mathbf{p})$ become unstable when the inner product $\langle\mathbf{w}_{r}\,,\mathbf{v}_{\theta}\rangle>0$. Thus, as a consequence of the Implicit Function Theorem and Proposition 6.1 of [15], this condition is equivalent to $\mathcal{C}>0$, where
@@
The parameter $\mathcal{C}$, defined in (13), characterises the direction of root crossing and, consequently, the type of Hopf bifurcation. Figure 2(b)-(c) visualises the delays $(\tau_{1},\tau_{2})$ categorised by this parameter. Red curves delineate regions where $\mathcal{C}<0$, while blue regions represent $\mathcal{C}>0$. Each intersection of a root of $q(\lambda,\boldsymbol{\tau},\mathbf{p})$ with the imaginary axis corresponds to a Hopf bifurcation. Specifically, a supercritical Hopf bifurcation occurs when the root crosses from left to right (i.e., $\mathcal{C}>0$), while a subcritical Hopf bifurcation occurs for a crossing from right to left (i.e., $\mathcal{C}<0$). The likelihood of these bifurcations increases with the difference between the delays, $|\tau_{1}-\tau_{2}|$. Moreover, the direction of root crossing exhibits a quasi-periodic behaviour, as illustrated in Figure 2(b)-(c). These results further suggest a dependence of the bifurcation type on the frequency. Specifically, the emergence of $\alpha$- and $\omega$-limit cycles is observed across a range of oscillatory frequencies. These limit cycles manifest at both lower frequencies (Figure 2(b)) and elevated frequencies (Figure 2(c)), a phenomenon associated with an alternating sequence of criticality transitions. This observation underscores the nuanced relationship between oscillatory dynamics and system criticality within the explored parameter set values.
@@
As established in the preceding section, the introduction of two delays significantly alters the system’s dynamics, inducing self-sustained oscillations through Hopf bifurcations. To further investigate this phenomenon, we fix the time delay $\tau_{1}$ at a specific value and slowly vary $\tau_{2}$. The resulting bifurcation diagram is shown in Figure 3, where their dynamical behaviour is initially depicted in Figure 2(b) and 2(c), where a countably infinite set of $\tau_{2}$ values at which Hopf bifurcations occur can be seen. This is illustrated by the distribution of asterisks along the dotted line $\tau_{1}=52$ in Figure 2(b) and 2(c), each marking a bifurcation point.
@@
Figure 3: Bifurcation diagram with respect to the delay parameter $\tau_{2}$, with $\tau_{1}=52$. A sample of countably set of Hopf bifurcations is observed, with ten representative bifurcations labelled $HB_{i}$, $i=1,\dots,10$. The point labeled $HB$, located between $HB_{6}$ and $HB_{7}$, corresponds to the first $HB$ point shown in Figure 2(b). Periodic orbit branches emanate from each $HB_{i}$ and $HB$ point. Stable branches consisted of either periodic or steady-state orbits are indicated in blue, while unstable branches are shown in red. Asterisks mark torus ($TR$) and fold ($LP$) bifurcation points detected along periodic orbit branches.
Figure 3 presents the Hopf bifurcations observed for a fixed delay $\tau_{1}=52$. A primary Hopf bifurcation, denoted $HB_{0}$, occurs at $\tau_{2}=10.4208$ within the high-frequency interval $\omega\in\Omega_{2}$. This primary bifurcation is followed by a sequence of secondary Hopf bifurcations, labeled $HB_{i}$, $i=1,\dots,10$. An additional Hopf bifurcation, $HB$, is observed at $\tau_{2}=118.0674$ for some $\omega\in\Omega_{1}$, as depicted in Figure 2(b). The remaining secondary $HB_{i}$ bifurcations also occur within the low- and high-frequency intervals $\Omega_{1}$ and $\Omega_{2}$, as can be seen in Figure 2(b) and 2(c). We pay special attention to the high-frequency interval $\Omega_{2}$. The bifurcation diagram in Figure 3 validates the computations presented in Figure 2. Specifically, the criticality of the bifurcations in Figure 3 is consistent with the predictions derived from (13), as indicated by the asterisks in Figure 2(c).
@@
Following the identification of the $\tau_{2}$ range associated with Hopf bifurcations, numerical continuation of equilibrium and periodic branches was performed using DDE-BIFTOOL [8]. Figure 3 displays these branches, including their stability properties (stable branches in blue, unstable branches in red), in a neighbourhood of the equilibrium point $E_{1}$, parameterised by $\tau_{2}\in[2,200]$ with $\tau_{1}=52$. The Hopf bifurcations marked with asterisks in Figure 2(c) were successfully located using DDE-BIFTOOL and correspond to the $HB_{i}$ points in Figure 3. Note that $HB_{0}$ is not shown in Figure 2(c) due to the figure’s scale. The $HB$ point in Figure 3 (between $HB_{6}$ and $HB_{7}$) corresponds to the first $HB\approx 117.3664$ in Figure 2(b). Slight numerical discrepancies were observed between the Hopf bifurcation points identified in Figures 2 and 3.
@@
Figure 4: Top Panel: two-parameter continuation of Hopf bifurcation branches in the ($b$, $\tau_{2}$) parameter plane, with $\tau_{2}\in[2,200]$. Panels (1)-(4): temporal dynamics of the system variables for selected parameter values: (1) $\tau_{2}=80$, $b=1.5$; (2) $\tau_{2}=105$, $b=1.36$; (3) $\tau_{2}=115$, $b=1.36$; and (4) $\tau_{2}=160$, $b=1.2$. Other parameter values as in Table 1.
Two-parameter continuation was then performed, varying $b$ and $\tau_{2}$ for each Hopf bifurcation point within the range $\tau_{2}\in[2,200]$. The parameter $b$, related to the decay rate of motile to static bacteria, was chosen as the secondary bifurcation parameter due to its critical role in influencing the interplay of self-sustained oscillatory dynamics, as discussed in [18]. The results of this analysis are presented in the top panel of Figure 4. The two-parameter continuation of the Hopf branches reveals the formation of intersecting islands as $\tau_{2}$ increases, suggesting the potential for complex oscillatory behaviour. To shed light on this possibility, specific values of $b$ and $\tau_{2}$ were selected to explore key system behaviours over time. These values are: $\tau_{2}=80$, $b=1.5$ (point 1); $\tau_{2}=105$, $b=1.36$ (point 2); $\tau_{2}=115$, $b=1.36$ (point 3); and $\tau_{2}=160$, $b=1.2$ (point 4). The corresponding time series are shown in panels 1-4 of Figure 4. While the temporal dynamics depicted in panels 1 through 3 are qualitatively similar, panel 4 suggests the emergence of complex, non-periodic oscillations. This behaviour may result from the combined effect of small $b$ values and large $\tau_{2}$ values, given that $\tau_{1}$ remains fixed at 52.
@@ 5 Intermittent chaos in time delay QS system
Figure 5: Upper row: Poincaré map computed on the cross-section $u=2$ (black dots) after transient time, as $\tau_{2}$ is slowly varied; Pearson correlation coefficient (blue bars) measuring the correlation between two initially nearby $w$-component orbits. Bottom row: frequency-amplitude plots and projections of representative orbits onto the ($u$, $v$) plane for three key distinct values of $\tau_{2}$; the central panel depicts a quasi-periodic orbit, while the left-hand and right-hand panels show stable periodic orbits. Other parameter values as in Table 1.
@@
To gain a comprehensive understanding of the oscillatory features governed by critical bifurcations, the Poincaré map of the time-delayed QS system (2) was computed over a substantial range of $\tau_{2}$ values, while holding $\tau_{1}$ constant. The first row of Figure 5 depicts the intersections of the state variable $u$ with the cross-section $u=2$ (black dots)—effectively, the Poincaré map—after transient time, as $\tau_{2}$ is slowly increased within the interval $\tau_{2}\in[2,1000]$ for fixed values of $\tau_{1}=52$ and $b=1.4$. As can be seen there, intervals of stable periodic solutions and complex quasi-periodic oscillations are interspersed until $\tau_{2}$ reaches a value sufficient to completely disrupt the periodic dynamics; see Figure 5, upper panel, $\tau_{2}\approx 700$. Beyond this point, the non-periodic oscillations suggest a chaotic regime, which is further explored in the bottom row of Figure 5 through frequency-amplitude plots and projections of distinguished orbits onto the $(u,v)$ phase plane. That is, for $\tau_{2}=250$, the Poincaré map indicates a stable periodic behaviour, which is corroborated by the well-defined orbit in the $(u,v)$ phase plane and the discrete frequency spectrum shown in the left-hand panel of the bottom row. When $\tau_{2}=380$, the dispersed Poincaré points and the dense orbit in the phase plane, coupled with the continuous frequency spectrum in the central panel, strongly indicate a chaotic regime. Finally, at $\tau_{2}=450$, the system returns to a stable periodic window, and the conclusions drawn for the case of $\tau_{2}=250$ apply.
@@
$r_{x,y}=\dfrac{\sum\limits_{i=1}^{n}(x_{i}-\overline{x})(y_{i}-\overline{y})}{\sqrt{\sum\limits_{i=1}^{n}(x_{i}-\overline{x})^{2}}\sqrt{\sum\limits_{i=1}^{n}(y_{i}-\overline{y})^{2}}}.$
@@
This coefficient ranges from -1 to 1, where 1 indicates perfect positive correlation, -1 perfect negative correlation, and 0 no correlation. In this context, the correlation was calculated between two discretised trajectories of the $w$ state-component, initialised with a difference of approximately $10^{-9}$ units. As shown in Figure 5, the blue bars reveal strong correlation between $w$-orbits within the periodic regimes, while the correlation collapses to naught in the windows where chaotic behaviour dominates. This characteristic pattern is indicative of intermittent chaos, a phenomenon first described by Pomeau and Manneville in their analysis of the Lorenz system [25]. Intermittent chaos is a well-established phenomenon in time-delayed regulatory gene circuits, often attributed to the inherent delays in intracellular processes [32].
@@
The comprehensive numerical simulations presented here provided further compelling evidence of the system’s richness dynamical repertoire. Such simulations not only corroborated the analytical predictions regarding the presence of self-sustained periodic oscillations, but also revealed the presence of intermittent chaotic behaviour as the delay parameters slowly vary. The intermittent nature of this chaotic regime, characterised by distinct phases of laminar flow punctuated by bursts of chaotic activity, strongly suggests a complex interaction between the destabilised steady-state and the emerging oscillatory modes. That is, the scheme that found and illustrated in Figure 5 corresponds to Type-II intermittency, within the Pomeau–Manneville classification design, which represents a distinct pathway to chaotic dynamics characterised by the irregular, intermittent switching between laminar phases and bursts of chaotic activity. A key distinguishing feature of Type-II intermittency is its association with a subcritical Hopf bifurcation. Specifically, as a system parameter traverses a critical value, a complex conjugate pair of eigenvalues associated with a stable periodic orbit crosses the imaginary axis, destabilising such orbit. The complex monodromic eigenvalues at the bifurcation point are of the form $\mu=\left(1+\varepsilon\right)e^{i\theta}$, where the local Poincare map can be described, in polar coordinates, by its normal form
$\left\{\begin{array}[]{l}r_{n+1}=\left(1+\varepsilon\right)r_{n}+ar^{3}_{n}\,,\\[4.30554pt] \theta_{n+1}=\theta_{n}+c+br^{2}\,,\end{array}\right.$
@@
where $a,b,c\in\mathbb{R}$ are constant, and Type-II intermittency occurs for $\varepsilon>0$, see [6]. This destabilisation gives rise not to a global chaotic attractor, but rather to a repeller in the vicinity of the now-unstable periodic orbit. The dynamics near this repeller are inherently chaotic, due to its unstable nature provided by positive Lyapunov exponents, here described by means of the geometric transversality parameter (13) and confirmed by computations shown in Figures 2 and 3. Orbits are intermittently “injected” into the neighbourhood of this repeller, leading to a burst of chaotic behaviour. Following this excursion into the chaotic region, the trajectory is subsequently re-injected into the phase space region previously occupied by the stable periodic orbit, resulting in a return to laminar periodic flow as is illustrated in Figure 5, upper row. The average durations of these laminar phases, $\bar{l}$, are statistically distributed, typically exhibiting a power-law scaling with the distance of the bifurcation parameter from its critical value, which is $\bar{l}\sim\varepsilon^{-1/2}$, see [6]. The essential mechanism underlying this phenomenon is the interplay between the destabilised periodic orbit (and its associated repeller) and the reinjection process that governs the transitions between laminar and chaotic phases.

### v3

@@ 2 Nonlinear time delayed Quorum Sensing system
Figure 2: The pseudo-characteristic polynomial (7) possesses at least one purely imaginary root when $\omega$ simultaneously satisfies the geometric conditions (11). This occurs within two distinct frequency intervals: a lower interval, $\Omega_{1}=[0,0.0466554]$, and a higher interval, $\Omega_{2}=[0.136335,0.187117]$, indicated by the grey bands in panel (a). Panels (b) and (c) illustrate all positive time delays for which the pseudo-polynomial roots lie on the imaginary axis, parameterised by $\omega$. Specifically, panel (b) shows trajectories for lower frequencies ($\omega\in\Omega_{1}$), while panel (c) displays those for higher frequencies ($\omega\in\Omega_{2}$). Asterisks mark the $(\tau_{1},\tau_{2})$ pairs where $\tau_{1}=52$. The direction in which roots cross the imaginary axis, as a function of time delay variation, is characterised by the transversality parameter (13): blue curves indicate a crossing from left to right on the imaginary axis, while red curves denote a crossing from right to left.
@@
Second, we consider the scenario where a pair of complex conjugate roots of $q(\lambda,\boldsymbol{\tau},\mathbf{p})$ crosses the imaginary axis into the right-hand side of the complex plane. In this case, the delay vector $\boldsymbol{\tau}=(\tau_{1},\tau_{2})$ shifts parallel to the vector $\mathbf{w}_{r}=(\partial\tau_{1}/\partial\sigma,\partial\tau_{2}/\partial\sigma)$. Consequently, if $\boldsymbol{\tau}\in\mathbb{R}_{+}^{2}$ lies to the left of a positively oriented orbit $\mathcal{T}^{k}$, two additional roots of $q(\lambda,\boldsymbol{\tau},\mathbf{p})$ become unstable when the inner product $\langle\mathbf{w}_{r}\,,\mathbf{v}_{\theta}\rangle>0$. Thus, as a consequence of the Implicit Function Theorem and Proposition 6.1 of [22], this condition is equivalent to $\mathcal{C}>0$, where
@@
The parameter $\mathcal{C}$, defined in (13), characterises the direction of root crossing and, consequently, the type of Hopf bifurcation. Figure 2(b)-(c) visualises the delays $(\tau_{1},\tau_{2})$ categorised by this parameter. Red curves delineate regions where $\mathcal{C}<0$, while blue regions represent $\mathcal{C}>0$. Each intersection of a root of $q(\lambda,\boldsymbol{\tau},\mathbf{p})$ with the imaginary axis corresponds to a Hopf bifurcation. Specifically, a supercritical Hopf bifurcation occurs when the root crosses from left to right (i.e., $\mathcal{C}>0$), while a subcritical Hopf bifurcation occurs for a crossing from right to left (i.e., $\mathcal{C}<0$). This leads directly to the following
Proposition 3.3.
Under conditions (11), the geometric transversality parameter (13) dictates the criticality of the Hopf bifurcation: (a) it is supercritical if $\mathcal{C}>0$, (b) and subcritical if $\mathcal{C}<0$.
The likelihood of these bifurcations increases with the difference between the delays, $|\tau_{1}-\tau_{2}|$. Moreover, the direction of root crossing exhibits a quasi-periodic behaviour, as illustrated in Figure 2(b)-(c). These results further suggest a dependence of the bifurcation type on the frequency. Specifically, the emergence of $\alpha$- and $\omega$-limit cycles is observed across a range of oscillatory frequencies. These limit cycles manifest at both lower frequencies (Figure 2(b)) and elevated frequencies (Figure 2(c)), a phenomenon associated with an alternating sequence of criticality transitions. This observation underscores the nuanced relationship between oscillatory dynamics and system criticality within the explored parameter set values.
@@
As established in the preceding section, the introduction of two delays significantly alters the system’s dynamics, inducing self-sustained oscillations through Hopf bifurcations. Such bifurcations associated with time delay combinations, specifically with fixed $\tau_{1}$ and varying $\tau_{2}$, are visually represented by the collection of asterisks along the dotted lines ($\tau_{1}=52$) in Figures 2(b) and 2(c). These figures demonstrate that both subcritical (red) and supercritical (blue) Hopf bifurcations occur in both lower ($\omega\in\Omega_{1}$) and higher ($\omega\in\Omega_{2}$) frequency ranges. To further investigate stability changes due to the introduction of time delays, we fixed the parameter $\tau_{1}$ as previously mentioned and slowly varied $\tau_{2}$. A bifurcation diagram is computed by using DDE-BIFTOOL [13] by extending a stability branch in the neighbourhood of equilibrium $E_{1}$ and numerically locating Hopf bifurcations within the range $\tau_{2}\in[2,200]$.
@@
Figure 3: Extended stability and periodic orbit continuation. The system’s stability and periodic orbit extensions with respect to the varying delay parameter $\tau_{2}$, with $\tau_{1}=52$ held fixed. The primary Hopf bifurcation $HB_{0}$ and ten secondary Hopf bifurcations ($HB_{i}$, for $i=1,\dots,10$) are identified, and their locations and criticality satisfactorily correspond with those predicted in Figure 2(c). The point labelled $HB$, situated between $HB_{6}$ and $HB_{7}$, corresponds to the first $HB$ point (not extended) shown in Figure 2(b). Periodic orbit branches emanate from each Hopf point. Stable branches, representing either periodic or steady-state orbits, are indicated in blue, while unstable branches are shown in red. Asterisks mark detected torus $TR$ and fold $LP$ bifurcation points along these periodic orbit branches.
The bifurcation diagram in Figure 3 validates the computations presented in Figure 2(b) and 2(c) regarding the location of Hopf points within the range $\tau_{2}\in[0,200]$. We observed slight numerical discrepancies between the identified Hopf bifurcation points in Figures 2 and 3, though these remained within acceptable numerical errors. Our analysis of criticality will now focus on the Hopf bifurcations shown in Figure 2(c), as the diversity of dynamics becomes increasingly complex in the higher frequency range $\omega\in\Omega_{2}$.
Specifically, the criticality of the bifurcations in Figure 2(c), predicted with the aid of the geometric transversality parameter (13), is consistent with the stability results presented in Figure 3. That is, $HB_{0}$ corresponds to a subcritical Hopf bifurcation, $HB_{1}$ to a supercritical Hopf bifurcation, and the remaining $HB_{i}$ for $i=2,...,10$ all denote subcritical Hopf bifurcations.
@@
Following the identification of the $\tau_{2}$ range associated with Hopf bifurcations, numerical continuation of equilibrium and periodic branches was performed using DDE-BIFTOOL. Figure 3 displays these branches, including their stability properties (stable branches in blue, unstable branches in red), in a neighbourhood of the equilibrium point $E_{1}$, parameterised by $\tau_{2}\in[2,200]$ with $\tau_{1}=52$. The Hopf bifurcations marked with asterisks in Figure 2(c) were successfully located using DDE-BIFTOOL and correspond to the $HB_{i}$ points in Figure 3. Note that $HB_{0}$ is not shown in Figure 2(c) due to the figure’s scale. The $HB$ point in Figure 3 (between $HB_{6}$ and $HB_{7}$) corresponds to the first $HB\approx 117.3664$ in Figure 2(b), which we did not perform a numerical continuation for this particular case, as our focus is directed towards the higher frequency regime characterised by $\omega\in\Omega_{2}$. Slight numerical discrepancies were observed between the Hopf bifurcation points identified in Figures 2 and 3.
@@
Figure 4: Top Panel: this panel presents a two-parameter continuation of Hopf bifurcations within the range $\tau_{2}\in[0,200]$, with $\tau_{1}=52$ fixed. The emergence of intersecting islands as $\tau_{2}$ increases suggests the onset of complex oscillations. Panels (1)-(4): to further investigate the complexity in the time evolution scenario, the temporal dynamics of the system variables $u(t)$, $v(t)$, and $w(t)$ (red, black, and blue lines, respectively) are computed for selected parameter values: (1) $\tau_{2}=80$, $b=1.5$; (2) $\tau_{2}=105$, $b=1.36$; (3) $\tau_{2}=115$, $b=1.36$; and (4) $\tau_{2}=160$, $b=1.2$. Other parameter values are as in Table 1. While panels (1) through (3) display qualitatively similar temporal dynamics, panel (4) indicates the onset of complex, non-periodic oscillations.
To further investigate the presence of complex oscillations on the system, a two-parameter continuation was then performed, varying $b$ and $\tau_{2}$ for each Hopf bifurcation point within the range $\tau_{2}\in[2,200]$. The parameter $b$, related to the decay rate of motile to static bacteria, was chosen as the secondary bifurcation parameter due to its critical role in influencing the interplay of self-sustained oscillatory dynamics, as discussed in [31]. The results of this analysis are presented in the top panel of Figure 4. The two-parameter continuation of the Hopf branches reveals the formation of intersecting islands as $\tau_{2}$ increases, suggesting the potential for complex oscillatory behaviour. To shed light on this possibility, specific values of $b$ and $\tau_{2}$ were selected to explore key system behaviours over time. These values are: $\tau_{2}=80$, $b=1.5$ (point 1); $\tau_{2}=105$, $b=1.36$ (point 2); $\tau_{2}=115$, $b=1.36$ (point 3); and $\tau_{2}=160$, $b=1.2$ (point 4). The corresponding time series are shown in panels 1-4 of Figure 4. While the temporal dynamics depicted in panels 1 through 3 are qualitatively similar, panel 4 suggests the emergence of complex, non-periodic oscillations. This behaviour may result from the combined effect of small $b$ values and large $\tau_{2}$ values, given that $\tau_{1}$ remains fixed at 52.
@@ 5 Intermittent chaos in time delay QS system
Figure 5: Upper row: Poincaré map computed on the cross-section $u=2$ (black dots) after transient time, as $\tau_{2}$ is slowly varied; Pearson correlation coefficient (blue bars) measuring the correlation between two initially nearby $w$-component orbits. Lower row: frequency-amplitude plots and projections of representative orbits onto the ($u$, $v$) plane for three key distinct values of $\tau_{2}$; the central panel depicts a quasi-periodic orbit, while the left-hand and right-hand panels show stable periodic orbits. Other parameter values as in Table 1.
@@
To gain a comprehensive understanding of the oscillatory features governed by critical bifurcations, the Poincaré map of the time-delayed QS system (2) was computed over a substantial range of $\tau_{2}$ values, while holding $\tau_{1}$ constant. The first row of Figure 5 depicts the intersections of the state variable $u$ with the cross-section $u=2$ (black dots)—effectively, the Poincaré map—after transient time, as $\tau_{2}$ is slowly increased within the interval $\tau_{2}\in[2,1000]$ for fixed values of $\tau_{1}=52$ and $b=1.4$. As can be seen there, intervals of stable periodic solutions and complex quasi-periodic oscillations are interspersed until $\tau_{2}$ reaches a value sufficient to completely disrupt the periodic dynamics; see Figure 5, upper panel, $\tau_{2}\approx 700$. Beyond this point, the non-periodic oscillations suggest a chaotic regime, which is further explored in the bottom row of Figure 5 through frequency-amplitude plots and projections of distinguished orbits onto the $(u,v)$ phase plane. That is, for $\tau_{2}=250$, the Poincaré map indicates a stable periodic behaviour as can be seen from the trend forming the black curve, which is corroborated by the well-defined orbit in the $(u,v)$ phase plane and the discrete frequency spectrum shown in the left-hand panel of the bottom row. When $\tau_{2}=380$, the dispersed Poincaré points and the dense orbit in the phase plane, coupled with the continuous frequency spectrum in the central panel, strongly indicate a chaotic regime. Finally, at $\tau_{2}=450$, the system returns to a stable periodic window, and the conclusions drawn for the case of $\tau_{2}=250$ apply.
Figure 6: Initial conditions sensitivity for varying delay parameter $\tau_{2}$. Upper panel: the largest Lyapunov exponent $\lambda_{\textrm{max}}$ is shown as a function of $\tau_{2}$. Positive values (blue shaded regions) indicate chaotic regimes, while values near zero denote self-sustained periodic oscillations. Lower panel: the mean curve separation $\mu\{\delta\}$ (initial perturbation $\delta_{0}\approx 10^{-7}$), in blue, illustrates divergence in the chaotic windows and convergence to naught in non-chaotic intervals after $t=17.5\times 10^{4}$. The accompanying Poincaré map of $u$ (black dots) in both panels visually confirms the system’s transitions between chaotic sets and stable limit cycles, displaying the intermittent behaviour.
@@
(14) $r_{x,y}=\dfrac{\sum\limits_{i=1}^{n}(x_{i}-\overline{x})(y_{i}-\overline{y})}{\sqrt{\sum\limits_{i=1}^{n}(x_{i}-\overline{x})^{2}}\sqrt{\sum\limits_{i=1}^{n}(y_{i}-\overline{y})^{2}}}.$
@@
This coefficient ranges from -1 to 1, where 1 indicates perfect positive correlation, -1 perfect negative correlation, and 0 no correlation. In this context, the correlation was calculated between two discretised orbits of the $w$ state-component, initialised with a difference of approximately $10^{-9}$ units. As shown in Figure 5, the blue bars reveal strong correlation between $w$-orbits within the periodic regimes, while the correlation collapses to naught in the windows where chaotic behaviour dominates. This characteristic pattern is indicative of intermittent chaos, a phenomenon first described by Pomeau and Manneville in their analysis of the Lorenz system [33]. Intermittent chaos is a well-established phenomenon in time-delayed regulatory gene circuits, often attributed to the inherent delays in intracellular processes [52].
To further analyse this phenomenon, we compute the largest Lyapunov exponent $\lambda_{\textrm{max}}$ using the method proposed by Rosenstein et al. [48], as implemented in the MATLAB routine lyapunovExponent [34]. The top panel of Figure 6 depicts $\lambda_{\textrm{max}}$ as a function of $\tau_{2}$. It is worth noting that $\lambda_{\textrm{max}}>0$ even in some seemingly non-chaotic regions. This observation may be a consequence of the phenomenon reported by Grantham et al. in [21] and Stefanski et al. in [51], where positive finite-time Lyapunov exponents (FTLEs) characterise the initial, transient behaviour of a system before it settles onto a stable limit cycle. Their work demonstrated an inverse relationship: larger positive FTLEs during the chaotic transient correspond to shorter “rambling times”, meaning the system escapes the chaotic phase more quickly to reach its periodic orbit. Conversely, smaller positive FTLEs indicate a more prolonged period of transient chaos before convergence to the limit cycle. This implies that even when a system ultimately approaches a stable non-chaotic attractor, FTLEs can quantify the intensity and persistence of its preceding chaotic transient. Moreover, it is important to recognise that numerically computed positive FTLEs can emerge as numerical artefacts of discretisation in systems theoretically exhibiting stable limit cycles. As demonstrated by De Markus et al. in [10], the integration step size can induce “computational chaos” where FTLEs become positive for periodic systems, or conversely, lead to numerically induced periodicity in truly chaotic ones. This highlights the crucial role of numerical methods in shaping observed dynamics, suggesting that positive FTLEs in stable systems often indicate numerical errors or transient chaos rather than sustained chaotic behaviour, for instance.
Figure 7: Non-chaotic regime for $\tau_{2}=100,280$. Upper panels: time series of state variables exhibit stable self-sustained oscillations for $x(t)$ and $x(t,\epsilon)$. Bottom left-hand panel: log-distance between perturbed trajectories. Bottom right-hand panel: the two phase-space orbits further confirms a stable limit cycle. Other parameter values as in Table 1.
To properly characterise our findings, $\lambda_{\textrm{max}}$ is presented as $\tau_{2}$-dependent alongside the mean trajectory separation distance $\mu\{\delta\}$ in the bottom panel of Figure 6. This quantity represents the average distance between two orbits after a sufficiently long time, specifically $t=17.5\times 10^{4}$, with an initial perturbation magnitude of $\delta_{0}\approx 10^{-7}$ between their initial conditions. To enable a comprehensive comparison of stable limit cycles and chaotic windows across both panels, a Poincaré map is superposed.
Observe that $\lambda_{\textrm{max}}$ clearly increases by nearly an order of magnitude as the system becomes chaotic. Conversely, it decreases when stable limit cycles are dominant, as illustrated in the top panel of Figure 6. This behaviour is further corroborated by $\mu\{\delta\}$, which attains significantly larger values in chaotic regions, consistent with exponential divergence, while dropping significantly during periodic oscillations, as shown in the bottom panel of Figure 6. The interplay between $\lambda_{\textrm{max}}$ and $\mu\{\delta\}$, combined with the corresponding Poincaré map, which exhibits no presence of accumulation points in chaotic regions, and the Pearson linear correlation coefficient given by equation (14), confirms the presence of intermittent chaos.
Figure 8: Chaotic regime for $\tau_{2}=178,350$. Top row: time series of the state variables exhibit quasi-periodic behaviour for $x(t)$ and $x(t,\epsilon)$. Bottom left- and right-hand panels: log-distance between perturbed trajectories, and the two phase-space orbits further show initial conditions sensitivity, respectively. Other parameter values as in Table 1.
To further illustrate these observations, Figures 7 and 8 depict the system’s orbits for selected values of $\tau_{2}$. Here, $\delta(x(t),x(t,\epsilon))$ denotes the Euclidean distance between a reference orbit $x(t):=(w(t),v(t),u(t))$ (initial state $x(0)$) and a perturbed orbit $x(t,\epsilon):=\left(w(t,\epsilon),v(t,\epsilon),u(t,\epsilon)\right)$ (initial condition $x(0)+\epsilon$, where $0<\|\epsilon\|\ll 1$). For $\tau_{2}=100,280$, where the Poincaré map indicates stable self-sustained oscillations, $\delta(x(t),x(t,\epsilon))$ remains bounded, confirming non-chaotic dynamics, as shown in Figure 7. In contrast, for $\tau_{2}=178,350$, $\delta(x(t),x(t,\epsilon))$ exhibits exponential-like growth during the transient phase, consistent with chaotic behaviour. Notice that, in non-chaotic regimes, $\delta(x(t),x(t,\epsilon))$ stabilises at significantly low values post-transient compared to chaotic cases where the stabilisation is not clearly observed, further validating the computation of an approximation of the largest Lyapunov exponent, as illustrated in Figure 8.
@@
The comprehensive numerical simulations presented here provided further compelling evidence of the system’s richness dynamical repertoire. Such simulations not only corroborated the analytical predictions regarding the presence of self-sustained periodic oscillations, but also revealed the presence of intermittent chaotic behaviour as the delay parameters slowly vary. The intermittent nature of this chaotic regime, characterised by distinct phases of laminar flow punctuated by bursts of chaotic activity, strongly suggests a complex interaction between the destabilised steady-state and the emerging oscillatory modes. That is, the scheme that found and illustrated in Figures 5 and 6 corresponds to Type-II intermittency, within the Pomeau–Manneville classification design, which represents a distinct pathway to chaotic dynamics characterised by the irregular, intermittent switching between laminar phases and bursts of chaotic activity. A key distinguishing feature of Type-II intermittency is its association with a subcritical Hopf bifurcation. Specifically, as a system parameter traverses a critical value, a complex conjugate pair of eigenvalues associated with a stable periodic orbit crosses the imaginary axis, destabilising such orbit. The complex monodromic eigenvalues at the bifurcation point are of the form $\mu=\left(1+{\rho}\right)e^{i\theta}$, where the local Poincare map can be described, in polar coordinates, by its normal form
$\left\{\begin{array}[]{l}r_{n+1}=\left(1+{\rho}\right)r_{n}+ar^{3}_{n}\,,\\[4.30554pt] \theta_{n+1}=\theta_{n}+c+br_{n}^{2}\,,\end{array}\right.$
@@
where $a,b,c\in\mathbb{R}$ are constant, and Type-II intermittency occurs for ${\rho>0}$, see [11]. This destabilisation gives rise not to a global chaotic attractor, but rather to a repeller in the vicinity of the now-unstable periodic orbit. The dynamics near this repeller are inherently chaotic, due to its unstable nature provided by positive Lyapunov exponents, here described by means of Proposition 3.3 and confirmed by computations shown in Figures 2, 35 and 6. Orbits are intermittently “injected” into the neighbourhood of this repeller, leading to a burst of chaotic behaviour. Following this excursion into the chaotic region, the orbit is subsequently re-injected into the phase space region previously occupied by the stable periodic orbit, resulting in a return to laminar periodic flow as is illustrated in Figure 5, upper row, and both panels of Figure 6. The average durations of these laminar phases, $\bar{l}$, are statistically distributed, typically exhibiting a power-law scaling with the distance of the bifurcation parameter from its critical value, which is $\bar{l}\sim{\rho}^{-1/2}$, see [11]. The essential mechanism underlying this phenomenon is the interplay between the destabilised periodic orbit (and its associated repeller) and the reinjection process that governs the transitions between laminar and chaotic phases.